### v1

# Resolvent4py: a parallel Python package for analysis, model reduction and control of large-scale linear systems

Abstract
In this paper, we present resolvent4py, a parallel Python package for the analysis, model reduction and control of large-scale linear systems with millions or billions of degrees of freedom. This package provides the user with a friendly Python-like experience (akin to that of well-established libraries such as numpy and scipy), while enabling MPI-based parallelism through mpi4py, petsc4py and slepc4py. In turn, this allows for the development of streamlined and efficient Python code that can be used to solve several problems in fluid mechanics, solid mechanics, graph theory, molecular dynamics and several other fields.
keywords:
Python , Parallel computing , Petsc4py , Slepc4py , Mpi4py , Model reduction , Resolvent analysis , Stability analysis , Harmonic resolvent analysis
journal: SoftwareX
Metadata
Table 1: Code metadata
Nr. Code metadata description Metadata
C1 Current code version v1.0.1
C2 Permanent link to code/repository used for this code version https://github.com/albertopadovan/resolvent4py
C3 Permanent link to Reproducible Capsule
C4 Legal Code License MIT
C5 Code versioning system used Git (GitHub)
C6 Software code languages, tools, and services used Python, Petsc4py, Slepc4py, Mpi4py
C7 Compilation requirements, operating environments & dependencies See README file in package distribution
C8 If available Link to developer documentation/manual https://albertopadovan.github.io/resolvent4py/
C9 Support email for questions alberto.padovan.94@gmail.com
1 Motivation and significance
The development of this package is motivated by recent (and not-so-recent) advances in the field of fluid mechanics, where linear analysis has been shown to provide great insights into the flow physics. For example, linear systems theory can be used to study the stability of fluid flows around equilibria, develop feedback control strategies to modify the flow behavior and compute reduced-order models to accelerate physical simulations. For systems of moderate dimension (e.g., fewer than 10,000 states), all these tasks can be performed straightforwardly with a few lines of Python code thanks to scientific computing libraries like scipy. For larger systems, the need for MPI-based parallelism significantly increases the implementation complexity of these algorithms, and it often requires the development of application-specific software written in compiled languages such as C, C++ and Fortran. The objective of this package is to provide a user-friendly pythonic environment similar to that of scipy and numpy, while exploiting the large-scale parallelism offered by the mpi4py, petsc4py and slepc4py libraries. In turn, this enables the development of user-friendly, streamlined and efficient code to perform day-to-day engineering tasks and physical analyses of large-scale systems.
Remark 1
Although the development of this package is motivated by problems in the field of fluid mechanics, resolvent4py is application agnostic, and its functionalities can be leveraged for any linear system of equations regardless of their origin.
2 Software description
resolvent4py is a Python library for the analysis, model reduction and control of large-scale linear systems. It relies on mpi4py [1] for multi-processing parallelism, and it leverages the data structures and functionalities provided by the petsc4py [2] and slepc4py [3] libraries.
2.1 Software architecture
At the core of this package is an abstract class, called LinearOperator, which serves as a blueprint for user-defined child classes that can be used to define a linear operator $L$. All child classes must implement at least three methods: apply(), which defines the action of $L$ on a vector, apply_mat(), which defines the action of $L$ on a matrix and destroy() to free the memory. The implementation of the second method may seem redundant to the reader who is familiar with native Python, where the dot() method can be used to apply the action of a linear operator to vectors and matrices alike. However, the need for this method is dictated by the fact that matrix-vector products and matrix-matrix products are performed with different functions and data structures in petsc4py.
If desired, additional methods can be implemented. For example, we often define the action of the hermitian transpose of a linear operator on vectors and matrices (apply_hermitian_transpose() and apply_hermitian_trans
pose_mat()
, respectively), the action of the inverse $L^{-1}$ on vectors, solve(), etcetera. Below is a list of concrete subclasses of LinearOperator that are provided by resolvent4py:
  1. 1.
    MatrixLinearOperator: the linear operator $L$ is defined by a sparse, MPI-distributed PETSc matrix.
  2. 2.
    LowRankLinearOperator: the linear operator $L$ is defined as the product of low-rank factors,
    $L=U\Sigma V^{*},$ (1)
    where $U$ and $V$ are tall and skinny matrices stored as MPI-distributed SLEPc BV structures [^1] The SLEPc BV structure (where BV stands for “Basis Vector”) holds a tall and skinny dense matrix whose rows are distributed across different MPI processors. The columns, however, are not distributed, and this allows for significantly reduced parallel overhead when the number of MPI processors is much larger than the number of columns, as is usually the case in practice. , while $\Sigma$ is a sequential two-dimensional numpy array of conformal size.
  3. 3.
    LowRankUpdatedLinearOperator: the linear operator $L$ is defined as $L=A+M$, where $A$ is itself a resolvent4py linear operator and $M$ is a low-rank linear operator, as described in the previous bullet point. An operator of this form arises in several applications and, if $A$ and $L$ are full-rank, then the inverse $L^{-1}$ can be computed using the Woodbury matrix inversion lemma [4],
    $L^{-1}=A^{-1}-A^{-1}U\Sigma\left(I+V^{*}A^{-1}U\Sigma\right)^{-1}V^{*}A^{-1}.$ (2)
  4. 4.
    ProjectionLinearOperator: the linear operator $L$ is defined either as $L=\Phi\left(\Psi^{*}\Phi\right)^{-1}\Psi^{*}$ or $L=I-\Phi\left(\Psi^{*}\Phi\right)^{-1}\Psi^{*}$, where $\Phi$ and $\Psi$ are tall and skinny matrices of size $N\times r$ stored as SLEPc BVs. In both cases, $L$ defines a projection (i.e., $L^{2}=L$).
  5. 5.
    ProductLinearOperator: the linear operator $L$ is defined as the product of other resolvent4py linear operators $L_{i}$ of conformal dimensions,
    $L=L_{m}L_{m-1}\ldots L_{2}L_{1}.$ (3)
2.2 Software functionalities
Once a linear operator is properly defined, resolvent4py currently allows for several types of analyses.
  1. 1.
    Eigendecomposition: the user can compute the left and right eigedecomposition of the linear operator $L$ using the Arnoldi algorithm [5]. The shift-and-invert technique can be used to compute the eigenvalues that are closest to a target value $s\in\mathbb{C}$, where $\mathbb{C}$ is the set of complex numbers.
  2. 2.
    Singular value decomposition (SVD): the user can compute an SVD of the linear operator using randomized linear algebra [6, 7]. This is useful for resolvent analysis [8, 9, 10] and harmonic resolvent analysis [11, 12, 13, 14]. For the specific case of resolvent analysis, the package also offers the possibility of computing the SVD using time-stepping techniques [15, 16].
  3. 3.
    Linear time-invariant balanced truncation: given the linear operator $L$, and appropriately-defined input and output matrices $B$ and $C$, the user may compute and balance the associated controllability and observability Gramians for model reduction purposes [17, 18, 19]. Specifically, we implement the algorithm outlined in [20].
Additionally, resolvent4py ships with several functions—available under the resolvent4py/utils directory and directly accessible to the user via the resolvent4py namespace—that further facilitate the use of our package and allow for streamlined application-specific code development. These include support for:
  1. 1.
    parallel I/O through petsc4py,
  2. 2.
    MPI-based communications using mpi4py,
  3. 3.
    manipulation of PETSc matrices/vectors and SLEPc BVs.
All these features are thoroughly documented using sphinx, and demonstrated with several examples that ship with the package.
2.3 Sample code snippet
In this subsection, we show a code snippet similar to the one used in section 3.2 to perform resolvent analysis on the hypersonic flow over a cone. The first thing performed by this piece of code is to read in the sparse complex-valued matrix $A$ of size $N\times N$. This matrix is stored on disk in sparse COO format by means of PETSc-compatible binary files "rows.dat", "columns.dat" and "values.dat". After assembling $A$ as a sparse PETSc matrix, we create a corresponding MatrixLinearOperator object and we perform resolvent analysis using the RSVD-$\Delta t$ algorithm presented in [15, 16]. The desired singular values and vectors of the resolvent operator are then saved to file using the parallel I/O routines available under the resolvent4py namespace. Running this piece of code with, e.g., $20$ processors can be done with the command mpiexec -n 20 python code_snippet.py.
Listing 1: Resolvent analysis via time stepping
import resolvent4py as res4py
# Define problem size and read jacobian from file
N = 121404 # Number of global rows (and columns) of A
Nloc = res4py.compute_local_size(N) # Number of local rows
file_names = ["rows.dat", "columns.dat", "values.dat"]
A = res4py.read_coo_matrix(file_names, ((Nloc, N), (Nloc, N)))
# Create resolvent4py linear operator
L = res4py.linear_operators.MatrixLinearOperator(A, ksp)
# Define the inputs to the rsvd-dt routine
dt = 1e-5 # Time step size
omega = 1.0 # Fundamental frequency (sets period T = 2pi/omega)
n_omegas = 5 # Number of harmonics of the fundamental to resolve
n_periods = 100 # Number of periods to integrate through
n_loop = 1 # Number of power iterations
n_rand = 30 # Number of random vectors
n_svals = 3 # Number of singular values/vectors to resolve
tol = 1e-3 # Tolerance to declare that transients have decayed
verbose = 1 # Verbosity level
Ulst, Slst, Vlst = res4py.linalg.resolvent_analysis_rsvd_dt(L, dt, omega, n_omegas, n_periods, n_loops, n_rand, n_svals, tol, verbose)
# Save to file
np.save("S.npy", Slst[1])
res4py.write_to_file("U.dat", Ulst[1])
res4py.write_to_file("V.dat", Vlst[1])
3 Illustrative examples
3.1 Hypersonic flow over a compression ramp
The first example is a Mach–$7.7$ flow over a $15\deg$ compression ramp of length $L=0.1\,m$. In the freestream, the Mach number is $M_{\infty}=7.7$, the pressure is $p_{\infty}=760\,Pa$, the temperature is $T_{\infty}=125\,K$, the streamwise velocity is $u_{\infty}=1726\,m/s$ and the Reynolds number based on $L$ is $Re_{\infty,L}=4.2\times 10^{5}$. This flow configuration is studied extensively in [21], and we refer to the latter for details on boundary conditions, spatial discretization, mesh and non-dimensionalization. As discussed in [21], this flow admits a two-dimensional steady state solution (see figure 1) that is unstable with respect to three-dimensional perturbations. These instabilities can be detected by linearizing the governing equations about the aforementioned steady state and studying the dynamics of three-dimensional, spanwise-periodic perturbations with wavenumber $\beta$. Upon linearization, the spatially-discretized compressible Navier-Stokes equations can be written compactly as
$\frac{d}{dt}q_{\beta}=A_{\beta}q_{\beta},\quad q_{\beta}\in\mathbb{C}^{N},$ (4)
where the state vector $q_{\beta}$ denotes the wavenumber-$\beta$ Fourier coefficients of the spatially-discretized flow variables. Here, given a computational grid of size $n_{y}\times n_{x}=236\times 1076$ and $5$ flow variables, the size of the state vector is $N=5\,n_{y}\,n_{x}\approx 1.27\times 10^{6}$. The eigenvalues of $A_{\beta}$ and corresponding eigenvectors for (non-dimensional) wavenumber $\beta=2\pi/0.066$ are computed using resolvent4py with $240$ processors on Stempede3 at Texas Advanced Computing Center. The eigenvalues are shown in figure 2, where we see good agreement with the results in Cao et al. [21]. Indeed, we find that the flow exhibits multiple instabilities corresponding to the five eigenvalues $\lambda$ with real part greater than zero. The flow structures associated with these instability modes are shown in figure 3, where we plot contours of the real part of the spanwise velocity component from the eigenvectors corresponding to three unstable eigenvalues. Here, we find that the eigenvector corresponding to $\lambda_{1}$ exhibits very good qualitative agreement with its counterpart in figure 14a in [21]. The eigenvectors corresponding to $\lambda_{2}$ and $\lambda_{3}$ are very similar in nature to those in figures 14b and 14c in [21], but it is important to observe that a direct comparison cannot be made since the modes shown in figures 14b and 14c in [21] correspond to wavenumbers $\beta=0.070$ and $\beta=0.079$, respectively.
Figure 1: Mach number contours of two-dimensional steady state solution for the Mach– $7.7$ flow over a $15\deg$ compression ramp.
Figure 2: Eigenvalues of the linearized Navier-Stokes equations at non-dimensional spanwise wavenumber $\beta=2\pi/0.066$ computed using resolvent4py (black dots). The red crosses denote the eigenvalues in [21] for comparison. The eigenvectors associated with $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are shown in figure 3.
Figure 3: Real part of the spanwise velocity component from the eigenvectors corresponding to eigenvalues $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ in figure 2.
3.2 Hypersonic flow over a cone
Another illustrative example is a Mach-5.9 flow over a 7 deg half-angle blunt cone. The freestream quantities are $M_{\infty}=5.9$, $p_{\infty}=3396.3$ $Pa$, $T_{\infty}=76.74$ $K$, $u_{\infty}=1036.1$ $m/s$, and $Re_{\infty}=18\times 10^{6}$ and the domain length is $L=0.4893\,m$. Bluntness effects produce unique instability phenomena in comparison with sharp cones, and are extensively studied using various DNS and operator theoretic methods in receptivity studies  [22]. At high Mach numbers, a discrete higher-order mode termed the second Mack mode becomes destabilized and dominates laminar-to-turbulent transition. This mechanism manifests itself as acoustic waves trapped within the boundary layer. Likewise, response modes are modulated by the entropy layer observed near the front of the cone, which can contribute to distinct turbulent transition mechanisms.
Figure 4: Mach number contours of two-dimensional steady state solution for the Mach–$5.9$ flow over a $7\deg$ blunt cone, truncated to solely the post-shock regime and normalized by domain length.
The spatially-discretized Navier-Stokes equations are linearized about the base flow shown in figure 4, and the singular value decomposition of the resolvent operator is computed using the RSVD-$\Delta$t algorithm using a script almost identical to the one shown in section 2.3. The temperature perturbations extracted from the leading right and left singular vectors of the resolvent are seen in figure  5. The response mode corresponding to maximal amplification appears trapped in the boundary layer, while the forcing mode follows the decaying entropy layer’s contour within the boundary layer, indicative of the importance of the high entropy gradients of the flow due to bluntness.
Figure 5: Temperature forcing and response modes computed using RSVD-$\Delta$t plotted in streamwise and wall-normal coordinates normalized by the streamwise length of the domain.
4 Impact
In the fluid mechanics community alone, linear systems tools like stability analysis [8, 21, 23], resolvent analysis [9, 10, 24], open-loop and feedback control design [25, 26, 27], model reduction of linear time-invariant and time-periodic systems via balanced truncation [17, 19, 20, 28], etc., are featured in hundreds of peer-reviewed publications and industry applications. However, despite the popularity of these methods and algorithms, there is no unified platform that helps with their streamlined implementation. resolvent4py aims to bridge this gap by offering a Python package that can significantly reduce the upfront cost associated with developing parallel code to answer pressing research questions in physics and engineering. Specifically, not only does resolvent4py ship with fully-parallel state-of-the-art algorithms [5, 6, 7, 20] that are used on a daily basis by researchers and engineers in several branches of physics, but it also offers a user-friendly, pythonic environment that allows for rapid code development, testing and deployment. In other words, our package can help scientists and engineers solve very large-scale problems of practical interest at a fraction of the human cost that would otherwise be necessary to develop application-specific software. The simple and self-contained code snippet in section 2.3 is a testament to the friendly environment enabled by our package.
Additionally, we believe that the current version of resolvent4py can serve as a stepping stone for additional software development—both by the original developers and the scientific community at large—that will be featured in future releases. For example, the infrastructure of resolvent4py enables the straightforward implementation of time-periodic balanced truncation using “Frequential Gramians” [28], harmonic resolvent analysis via time stepping [29], wavelet-based resolvent analysis [30] and One-Way Navier-Stokes (OWNS) for slowly-developing flows [31].
5 Conclusions
In this manuscript, we introduced resolvent4py, a parallel Python package for the analysis, model reduction and control of large-scale linear systems. Large-scale linear systems with millions or billions of degrees of freedom are ubiquitous in the sciences and engineering, and this package is built to provide a friendly pythonic environment for rapid development and deployment of MPI-based parallel software to solve problems of practical interest. Our library leverages mpi4py for MPI-based parallelism, and it uses the functionalities and data structures provided by petsc4py and slepc4py to enable stability analysis, input/output analysis, model reduction, control and optimization of linear time-invariant and time-periodic systems. Additionally, the package ships with several functions and features that are available to the user through the resolvent4py namespace and that facilitate user-specific and application-specific software development within the larger resolvent4py infrastructure.
Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. 2139536, issued to the University of Illinois at Urbana-Champaign by the Texas Advanced Computing Center under subaward UTAUS-SUB00000545 with Dr. Daniel Stanzione as the PI. DB gratefully acknowledges support from the Office of Naval Research (N00014-21-1-2256), and CR gratefully acknowledges support from the Air Force Office of Scientific Research (FA9550-19-1-0005). TC and VA gratefully acknowledge support from The Boeing Company (CT-BA-GTA-1) and the Office of Naval Research (N00014-25-1-2072). BV was supported by the LDRD Program at Sandia National Laboratories. Sandia is managed and operated by NTESS under DOE NNSA contract DE-NA0003525. The computations in section 3.1 were performed on TACC’s Stampede3 under under ACCESS allocation CTS090004.